Q4: Using souce transformation, find i in the circuit of Fig. 4. 1 30 5 mH w 50 8 sin 2001+ 30) A 1 m Solution:
Answer #1

Given 32 SmH Marm mF 52 @s sin (2005 +30) using Source transformation V= 8 Sin (2006 +38) * 5 V = 40 sin (2005 +38) = 40/30°